OpenStudy (anonymous):

PLEASE HELP QUICK I HAVE 16 story problem questions?????? PLEASE ANYONE SMART HELP???????

OpenStudy (anonymous):

what are they?

OpenStudy (anonymous):

We won't just give you the answers, but we can help you learn how to solve them. I doubt that any of us will make it through all 16 though!

OpenStudy (anonymous):

18.) A quality control inspecto has fund that 3.2% of the garmets produced at Stanford garments contain a defect. If standard garments produces 4117 garments in one day how many of those garments are expected to have a defect????????????

OpenStudy (anonymous):

Do you remember how to find the % of a number?

OpenStudy (anonymous):

ok, to find a percentage of something, there are several ways to find it, the easiest is to multiply the percent as a decimal to the number

OpenStudy (anonymous):

times by 100?

OpenStudy (anonymous):

i give up...good luck!

OpenStudy (texaschic101):

3.2% of 4117 turn the percent into a decimal " of " means multiply so once you get your decimal, multiply it by 4117

OpenStudy (anonymous):

no, to make a percent into a decimal, move the decimal place 2 places to the left. so what would 3.2% be as a decimal?

OpenStudy (anonymous):

.032 x4117???? then after i do that i get the nswer to the problem??? just asking

OpenStudy (texaschic101):

you can either move it two spaces to the left, or divide it by 100

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

i asked is that what i said what i do?

OpenStudy (texaschic101):

3.2 divided by 100 is what ?

OpenStudy (anonymous):

oh my goodness, @texaschic101 , i know you are trying to help, but hold on one second, moving the decimal place over 2 places is the same thing

OpenStudy (anonymous):

so @caitlynmcdaniel .032 x 4117 is what you need to do

OpenStudy (anonymous):

131.744 is what i got... am i done w/this problem??

OpenStudy (anonymous):

that would be how many have defects out of the 4117 products, so yes, that is the end of the problem

OpenStudy (texaschic101):

you could probably round it to 132

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

19.)Lantanya made a 17-minute phone call from her hotel in france and was charged 17 euro. At the time, $1 was worth 0.82 euro. Find the cost per minte of he call in dollars????????

OpenStudy (anonymous):

so for this one, you have to find how many times .82 goes into 17

OpenStudy (anonymous):

20.7 is what i got when i divided those two numbers

OpenStudy (anonymous):

am i done with that problem already?

OpenStudy (anonymous):

good, so that is how many dollars it cost.

OpenStudy (anonymous):

almost, now divide 20.7 by 17

OpenStudy (anonymous):

because you still have to find the cost per minute, you have the total cost of the call

OpenStudy (anonymous):

i got 1.22??

OpenStudy (anonymous):

and yes, that is the cost per minute

OpenStudy (anonymous):

1.22 answer? or now what??

OpenStudy (anonymous):

thats the answer

OpenStudy (anonymous):

thanks! 20.) ABC has vertices A(0,0) B(6,0) C(6,-4.5). triangle ADE is similar to ABC with a vertex at D (8,0). graph ABC and ADE on the same grid??????????????

OpenStudy (anonymous):

alright, so you need to draw a grid and draw triangle ABC

OpenStudy (anonymous):

doing now.. pleaseeee don't leave

OpenStudy (anonymous):

i wont, i honestly have done all my homework, and i have nothing better to do, so ill be here for a while

OpenStudy (anonymous):

got it. finished that one.. lets try 21.. And i APPRECIATE THE HELP (:

OpenStudy (anonymous):

no problem!

OpenStudy (anonymous):

21.) A basketball rim 10 ft high casts a shadow 15 ft. long. At the same time, a narby building casts a shadow that is 54 ft. long.. How tall is the building??????????

OpenStudy (anonymous):

yeesh, let me think for a second, i havent done one of these for 2 years lol.

OpenStudy (anonymous):

@jim_thompson5910 , could you help me out with this last one? i seem to have forgotten how to do it

jimthompson5910 (jim_thompson5910):

10/15 = x/54 10*54 = 15x 15x = 540 x = ???

OpenStudy (anonymous):

ok @caitlynmcdaniel , you got that? lol

OpenStudy (anonymous):

yes thanks!

OpenStudy (anonymous):

haha youre welcome!

OpenStudy (anonymous):

8.) Dan works as a house painter. He knows that it is safe to place the base of his 10 foot ladder 3 feet from a base of a house. Today h has to use a 25 foot ladder. Dan wants to keep the same ratios in order to be safe. HOW far should Dan place the base of his 25-foot ladder from the base of his house???????????????

OpenStudy (anonymous):

its similar to the previous one, do you think you can set it up?

OpenStudy (anonymous):

no/:

jimthompson5910 (jim_thompson5910):

solve for x 10/3 = 25/x

OpenStudy (anonymous):

cross multiply????

jimthompson5910 (jim_thompson5910):

yes, giving you what?

OpenStudy (anonymous):

75=10x you divide by 10.. i got x=7.5 for my answer?????

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

9.) the school newspaper took a survey. of the students polled, 15% said they did not have too much homework. sixty students were polled for the survey. how many students said they did not have toooo much homework???????????

jimthompson5910 (jim_thompson5910):

part/whole = percent/100 x/60 = 15/100 x = ??

OpenStudy (anonymous):

ok, remember the first problem we did?

OpenStudy (anonymous):

thanks @seesawn but i think jim Thompson already helped w/this one... thanks soooo much both of u!!!! (:

OpenStudy (anonymous):

ok, and no problem!

OpenStudy (anonymous):

i got 9 studens?????

OpenStudy (anonymous):

thats what i got as well

OpenStudy (anonymous):

10.) Cheryl wants to measure the distance across a stream. She took some measurements and drew a diagram. how wide is the stream??????

OpenStudy (anonymous):

DIAGRAM: 1 acute triangle pointing straight down... 20yd on top.. one acute triangle touching tips pointing straight up 15 yds bottom 17 yds right side????????????

OpenStudy (anonymous):

8 problems i need help with left!?!?!? (: you two have been sooooo much help!!!! (:

OpenStudy (anonymous):

could you draw it? or attach the image if there is one?

OpenStudy (anonymous):